what is the solution to the equation 2x+5=6x-5?
Found 2 solutions by FrankM, stanbon:
Answer by FrankM(1040)

2x+5=6x-5
Add 5 to both sides, subtract 2x from both sides
10=4x
x = 2.5

Answer by stanbon(75887)

what is the solution to the equation
2x+5=6x-5
-------
4x = 10
x = 5/2
x = 2 1/2
=================
Cheers,
Stan H.
=================